\begin{thebibliography}{99}

%Zur Quantenmechanik des magnetischen Elektrons
\bibitem{1} W. Pauli, \href{https://doi.org/10.1007/BF01397326}{Z. Phys. \textbf{43}, 601 (1927).}

% The Effect of Retardation on the Interaction of Two Electrons
\bibitem{2} G. Breit, \href{https://doi.org/10.1103/PhysRev.34.553}{Phys. Rev. \textbf{34}, 553 (1929).}

% On the Dirac theory of spin 1/2 particle and its nonrelativistic limit
\bibitem{3} L. L. Foldy, S. A, Wouthuysen, \href{https://doi.org/10.1103/PhysRev.78.29}{Phys. Rev. \textbf{78}, 29 (1950).}

% Intrinsic Magnetic Moment as a Nonrelativistic Phenomenon
\bibitem{4} A. Galindo, C. Sanchez del Rio, \href{https://doi.org/10.1119/1.1937856}{Am. J. Phys. \textbf{29}, 582 (1961).}

% On the Pauli–Schrödinger equation
\bibitem{5} T. C. Chapman, O. Cerceau, \href{https://doi.org/10.1119/1.13914}{Am. J. Phys. \textbf{52}, 994 (1984).}

% The quantum mechanical current of the Pauli equation
\bibitem{6} M. Nowakowski, \href{https://doi.org/10.1119/1.19149}{Am. J. Phys. \textbf{67}, 916 (1999).}

% Foldy-Wouthuysen transformation for a Dirac-Pauli dyon and the Thomas-Bargmann-Michel-Telegdi equation
\bibitem{7} T. -W. Chen, D. -W. Chiou, \href{https://doi.org/10.1103/PhysRevA.82.012115}{Phys. Rev. A \textbf{82}, 012115 (2010).}

%The Stationary Dirac Equation as a Generalized Pauli Equation for Two Quasiparticles
\bibitem{8} N. L. Chuprikov, \href{https://doi.org/10.1007/s10701-015-9888-3}{Found. Phys. \textbf{45}, 644 (2015).} 

% The Pauli and Lévy-Leblond equations, and the spin current density
\bibitem{9} J. M. Wilkes, \href{https://doi.org/10.1088/1361-6404/ab7495}{Eur. J. Phys. \textbf{41}, 035402 (2020).}


\bibitem{10} V. I. Fushchich, A. G. Nikitin. \emph{Symmetries of equations of quantum mechanics.} (New York: Allerton Press, 1994).

\bibitem{11} F. Gieres (Ed.). \emph{Symmetries in Physics.} (Atlantica S\'{e}guier Fronti\`{e}res,  1997).

	
% Symmetries of the Schrödinger–Pauli equation for neutral particles
\bibitem{12} A. G. Nikitin, \href{https://doi.org/10.1063/5.0021725}{J. Math. Phys. \textbf{62}, 083509 (2021).}


% Symmetries of Schrödinger–Pauli equations for charged particles and quasirelativistic Schrödinger equations
\bibitem{13} A. G. Nikitin, \href{https://doi.org/10.1088/1751-8121/ac515d}{ J. Phys. A: Math. Theor. \textbf{55}, 115202 (2022).}


% The effect of the spin-orbit interaction on the electronic structure of magnetic materials
\bibitem{14} B. I. Min, Y. -R. Jang, \href{https://doi.org/10.1088/0953-8984/3/27/006}{J. Phys.: Condens. Matter \textbf{3}, 5131 (1991).}



% The Zeeman effect for the 'anistropic hydrogen atom' in the complex WKB approximation: I. Quantization of closed orbits for the Pauli operator with spin-orbit interaction
\bibitem{15} V. V. Belov, V. M. Oliv\'{e}, J. L. Volkova, \href{https://doi.org/10.1088/0305-4470/28/20/013}{J. Phys. A: Math. Gen. \textbf{28}, 5799 (1995).}

% The Zeeman effect for the 'anisotropic hydrogen atoms' in the complex WKB approximation: II. Quantization of two-dimensional Lagrangian tori (with focal points) for the Pauli operator with spin-orbit interaction
\bibitem{16} V. V. Belov, V. M. Oliv\'{e}, J. L. Volkova, \href{https://doi.org/10.1088/0305-4470%2F28%2F20%2F014}{J. Phys. A: Math. Gen. \textbf{28}, 5811 (1995).}


% Quantum theory as a description of robust experiments: Derivation of the Pauli equation
\bibitem{17} H. De Raedt, M. I. Katsnelson, H. C. Donker, K. Michielsen, \href{https://doi.org/10.1016/j.aop.2015.04.017}{Ann. Phys. \textbf{359}, 166 (2015).}

% Sufficient condition for the openness of a many-electron quantum system from the violation of a generalized Pauli exclusion principle
\bibitem{18} R. Chakraborty, D. A. Mazziotti, \href{https://doi.org/10.1103/PhysRevA.91.010101}{Phys. Rev. A \textbf{91}, 010101 (2015).}

% Role of the electromagnetic momentum in the spin-orbit interaction
\bibitem{19} G. Spavieri, \href{https://doi.org/10.1140/epjd/e2016-60575-1}{Eur. Phys. J. D \textbf{70}, 263 (2016).}

% Quantum-mechanical force balance between multipolar dispersion and Pauli repulsion in atomic van der Waals dimers
\bibitem{20} O. Vaccarelli, D. V. Fedorov, M. Stöhr, A. Tkatchenko,  \href{https://doi.org/10.1103/PhysRevResearch.3.033181}{Phys. Rev. Research \textbf{3}, 033181 (2021).}

%
%\bibitem{21} A. G. Nikitin, "Symmetries of the Schroedinger-Pauli equations for charged particles and quasirelativistic Schroedinger equations." arXiv preprint arXiv:2109.08528 (2021).

% Study of the most relevant spin–orbit coupling descriptions of magnetic excitations in a series of lanthanide complexes
\bibitem{22} S. V. Rao, M. Piccardo, A. Soncini, \href{https://doi.org/10.1039/D1CP05469D}{Phys. Chem. Chem.
Phys. \textbf{24}, 9007 (2022).}

% Bohr's 1913 molecular model revisited
\bibitem{23} A. A. Svidzinsky, M. O. Scully,  D. R. Herschbach, \href{https://doi.org/10.1073/pnas.0505778102}{PNAS \textbf{102}, 11985 (2005).}

% Rutherford–Bohr Model of the Atom.
\bibitem{24} B. E. Podgo\u{s}ak, \href{https://doi.org/10.1007/978-3-319-25382-4}{\emph{Rutherford–Bohr Model of the Atom.}} Ch:3 p:143-180 in {Radiation Physics for Medical Physicists}, (Springer: E-book 2016).

% Spin-forbidden chemistry within the Breit-Pauli approximation
\bibitem{25} D. R. Yarkony, \href{https://doi.org/10.1080/01442359209353270}{Int. Rev. Phys. Chem. \textbf{11}, 195 (1992).}

% The Pauli potential from the differential virial theorem
\bibitem{26} A. Nagy, \href{https://doi.org/10.1002/qua.22497}{Int. J. Quantum Chem. \textbf{110}, 2117 (2010).} 

% Introduction to relativistic quantum chemistry.
\bibitem{27} K. G. Dyall,  K. F\ae gri Jr. \href{https://doi.org/10.1093/oso/9780195140866.001.0001}{\emph{Introduction to relativistic quantum chemistry.}} (New York: Oxford University Press, 2007).

% A Full Quantum Eigensolver for Quantum Chemistry Simulations
\bibitem{28} S. Wei, H. Li, G. Long, \href{https://doi.org/10.34133/2020/1486935}{Research \textbf{2020}, 1486935 (2020).} 


%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

% Differential-difference operators associated to reflection groups
\bibitem{Dunkl1989} C. F. Dunkl,  \href{https://doi.org/10.1090/S0002-9947-1989-0951883-8}{T. Am. Math. Soc. \textbf{311}, 167 (1989).}

% Reflection groups in analysis and applications
\bibitem{Dunkl2008} C. F. Dunkl, \href{https://doi.org/10.1007/s11537-008-0819-3}{Jpn J. Math. \textbf{3}, 215 (2008).}


% Encyclopedia of Mathematics and its Applications : Orthogonal Polynomials of Several Variables
\bibitem{Dunkl2014} C. F. Dunkl, Y. Xu, \emph{Encyclopedia of Mathematics and its Applications: Orthogonal Polynomials of Several Variables} (Cambridge: Cambridge University Press, 2014).




% Common algebraic structure for the Calogero - Sutherland models
\bibitem{Kakei1996} S. Kakei, \href{https://doi.org/10.1088/0305-4470/29/24/002}{ J. Phys. A: Math. Gen. \textbf{29}, L619 (1996).}

% Exact operator solution of the Calogero-Sutherland model
\bibitem{Lapointe1996} L. Lapointe, L. Vinet, \href{https://doi.org/10.1007/BF02099456}{Commun. Math. Phys. \textbf{178}, 425 (1996).} 

% Dunkl Operator Formalism for Quantum Many-Body Problems Associated with Classical Root Systems
\bibitem{Hikami1996} K. Hikami, \href{https://doi.org/10.1143/JPSJ.65.394}{J. Phys. Soc. Jpn. \textbf{65}, 394 (1996)}


% A Note on the Quantum Rule of the Harmonic Oscillator
\bibitem{Yang1951} L. M. Yang, \href{https://doi.org/10.1103/PhysRev.84.788}{Phys. Rev. \textbf{84}, 788 (1951).}

% Do the Equations of Motion Determine the Quantum Mechanical Commutation Relations?
\bibitem{Wigner1950} E. P. Wigner, \href{https://doi.org/10.1103/PhysRev.77.711}{Phys. Rev. \textbf{77}, 711 (1950).}

% A Generalized Method of Field Quantization
\bibitem{Green1953} H. S. Green, \href{https://doi.org/10.1103/PhysRev.90.270}{ Phys. Rev. \textbf{90}, 270 (1953).}

% Spin and Unitary-Spin Independence in a Paraquark Model of Baryons and Mesons
\bibitem{Greenberg1964} O. W. Greenberg, \href{https://doi.org/10.1103/PhysRevLett.13.598}{Phys. Rev. Lett. \textbf{13}, 598 (1964). }

% Deformed Heisenberg algebra and fractional spin field in 2 + 1 dimensions
\bibitem{Plyushchay1994} M. S. Plyushchay. \href{https://doi.org/10.1016/0370-2693(94)90828-1}{Phys. Lett. B \textbf{320}, 91 (1994).}  

% Deformed Heisenberg Algebra, Fractional Spin Fields, and Supersymmetry without Fermions
\bibitem{Plyushchay1996} M. S.  Plyushchay,  \href{https://doi.org/10.1006/aphy.1996.0012}{Ann. Phys. \textbf{245}, 339 (1996).} 



% Deformed Heisenberg algebra with reflection
\bibitem{Plyushchay1997} M. S. Plyushchay, \href{https://doi.org/10.1016/S0550-3213(97)00065-5}{Nucl. Phys. B \textbf{491}, 619 (1997).} 


%Three aspects of bosonized supersymmetry and linear differential field equation with reflection
\bibitem{Gamboa1999} J. Gamboa, M. S.  Plyushchay, J. Zanelli, \href{https://doi.org/10.1016/S0550-3213(98)00832-3}{Nucl. Phys. B \textbf{543}, 447 (1999).} 


% HIDDEN NONLINEAR SUPERSYMMETRIES IN PURE PARABOSONIC SYSTEMS
\bibitem{Plyushchay2000} M. Plyushchay, \href{https://doi.org/10.1142/S0217751X00001981}{Int. J. Mod. Phys. A \textbf{15}, 3679 (2000).}

% Fractional helicity, Lorentz symmetry breaking, compactification and anyons
\bibitem{Klishevich2001} S. M. Klishevich, M. Plyushchay, M. Rausch de Traubenberg, \href{https://doi.org/10.1016/S0550-3213%2801%2900442-4}{Nucl. Phys. B \textbf{616}, 419 (2001).}

% Anyon wave equations and the noncommutative plane
\bibitem{Horvathy2004} P. A. Horvathy, M. Plyushchay, \href{https://doi.org/10.1016/j.physletb.2004.05.043}{Phys. Lett. B \textbf{595}, 547 (2004).} 


% On the hydrogen atom via the Wigner–Heisenberg algebra
\bibitem{Rodrigues2009} R. de Lima Rodrigues, \href{https://doi.org/10.1088/1751-8113/42/35/355213}{J. Phys. A \textbf{42}, 355213 (2009).}

% Bosons, fermions and anyons in the plane, and supersymmetry
\bibitem{Horvathy2010} P. A. Hortv\'athy, M. Plyushchay, M. Valenzuela, \href{https://doi.org/10.1016/j.aop.2010.02.007}{Ann. Phys. \textbf{325}, 1931 (2010).}

% Thermodynamics of boson systems related to Dunkl differential–difference operators
\bibitem{Ubriaco2014} M. R. Ubriaco, \href{https://doi.org/10.1016/j.physa.2014.06.087}{Physica A \textbf{414}, 128 (2014).}


% Dunkl-supersymmetric orthogonal functions associated with classical orthogonal polynomials 
\bibitem{Luo2020} Y. Luo, S. T. Sujimoto, L. Vinet, A. Zhedanov, \href{https://doi.org/10.1088/1751-8121/ab63a9}{J. Phys. A: Math. Theor. \textbf{53}, 085205 (2020).}


%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%5

% The Dunkl oscillator in the plane: I. Superintegrability, separated wavefunctions and overlap coefficients
\bibitem{Genest20131} V. X. Genest, M. E. H. Ismail, L. Vinet, A. Zhedanov, \href{https://doi.org/10.1088/1751-8113/46/14/145201}{J. Phys. A: Math. Theor. \textbf{46}, 145201 (2013).}

% The singular and the 2:1 anisotropic Dunkl oscillators in the plane
\bibitem{Genest20132} V. X. Genest, L. Vinet, A. Zhedanov, \href{https://doi.org/10.1088/1751-8113/46/32/325201}{J. Phys. A: Math. Theor. \textbf{46}, 325201 (2013).} 

% The Hahn superalgebra and supersymmetric Dunkl oscillator models
\bibitem{Genest20133} V. X. Genest, J. M. Lemay, L. Vinet, A. Zhedanov, \href{https://doi.org/10.1088/1751-8113/46/50/505204}{J. Phys. A: Math. Theor. \textbf{46}, 505204 (2013).} 

% The Dunkl Oscillator in the Plane II: Representations of the Symmetry Algebra
\bibitem{Genest20141} V. X. Genest, M. E. H. Ismail, L. Vinet, A. Zhedanov, \href{https://doi.org/10.1007/s00220-014-1915-2}{Commun. Math. Phys. \textbf{329}, 999 (2014).}

% The Dunkl oscillator in three dimensions
\bibitem{Genest20142} V. X. Genest L. Vinet, A. Zhedanov, \href{https://doi.org/10.1088/1742-6596/512/1/012010}{J. Phys.: Conf. Ser. \textbf{512}, 012010 (2014).}

% The Dunkl–Coulomb problem in the plane
\bibitem{Vincent3} V. X. Genest, A. Lapointe, L. Vinet, \href{https://doi.org/10.1016/j.physleta.2015.01.023}{Phys. Lett. A \textbf{379}, 923 (2015).}

% Families of 2D superintegrable anisotropic Dunkl oscillators and algebraic derivation of their spectrum
\bibitem{Isaac2016} P. S. Isaac, I. Marquette, \href{https://doi.org/10.1088/1751-8113/49/11/115201}{J. Phys. A: Math. Theor. \textbf{49}, 115201 (2016).}

% SU(1,1) solution for the Dunkl oscillator in two dimensions and its coherent states
\bibitem{Salazar2017} M. Salazar-Ram\'irez, D. Ojeda-Guill\'en, R. D. Mota, V. D. Granados, \href{https://doi.org/10.1140/epjp/i2017-11314-3}{Eur. Phys. J. Plus \textbf{132}, 39 (2017).}

% SU(1,1) solution for the Dunkl–Coulomb problem in two dimensions and its coherent states
\bibitem{Salazar2018} M. Salazar-Ram\'irez, D. Ojeda-Guill\'en, R. D. Mota, V. D. Granados, \href{https://doi.org/10.1142/S0217732318501122}{Mod. Phys. Lett. A \textbf{33}, 1850112 (2018).}

% Exact solution of the relativistic Dunkl oscillator in (2+1) dimensions
\bibitem{Mota2019} R. D. Mota, D. Ojeda-Guillén, M. Salazar-Ramírez, V. D. Granados, \href{https://doi.org/10.1016/j.aop.2019.167964}{Ann. Phys. \textbf{411}, 167964 (2019).}

% The Dunkl–Coulomb problem in three-dimensions: energy spectrum, wave functions and h-spherical harmonics
\bibitem{Ghazouani2019} S. Ghazouani, I. Sboui, M. A. Amdouni, M. B. El Hadj Rhouma, \href{https://doi.org/10.1088/1751-8121/ab0d98}{ J. Phys. A: Math. Theor. \textbf{52}, 225202 (2019).}


% Superintegrability of the Dunkl–Coulomb problem in three-dimensions
\bibitem{Ghazouani2020} S. Ghazouani, S. Insaf, \href{https://doi.org/10.1088/1751-8121/ab4a2d}{J. Phys. A: Math. Theor. \textbf{53}, 035202 (2020).} 

% Algebraic approach to the Dunkl–Coulomb problem and Dunkl oscillator in arbitrary dimensions
\bibitem{Ghazouani2021} S. Ghazouani, \href{https://doi.org/10.1007/s13324-020-00470-4}{Anal. Math. Phys. \textbf{11}, 35 (2021).}

% Thermal Properties of Relativistic Dunkl Oscillators
\bibitem{Hamil20221} B. Hamil, B. C. L\"{u}tf\"{u}o\u{g}lu, \href{https://doi.org/10.1140/epjp/s13360-022-03055-1}{Eur. Phys. J. Plus \textbf{137}, 812 (2022).}

% Dunkl graphene in constant magnetic field
\bibitem{Hamil20222} B. Hamil, B. C. L\"{u}tf\"{u}o\u{g}lu, \href{https://doi.org/10.1140/epjp/s13360-022-03463-3}{Eur. Phys. J. Plus \textbf{137}, 1241 (2022).}

% Dunkl–Klein–Gordon Equation in Three-Dimensions: The Klein–Gordon Oscillator and Coulomb Potential
\bibitem{Hamil20223} B. Hamil, B. C. L\"{u}tf\"{u}o\u{g}lu, \href{https://doi.org/10.1007/s00601-022-01776-8}{Few-Body Syst. \textbf{63}, 74 (2022).}

% Bound states of the isotonic Mathews–Lakshmanan oscillator system within the Dunkl formalism
\bibitem{43} A. Schulze-Halberg, \href{https://doi.org/10.1142/S0217732322501784}{Mod. Phys. Lett. A \textbf{37}, 2250178 (2022).}

% Relativistic solutions of generalized-Dunkl Harmonic and Anharmonic Oscillators 
\bibitem{Samira} S. Hassanabadi, J. K\u{r}\'i\u{z}, B.C. L\"{u}tf\"{u}o\u{g}lu, H. Hassanabadi, \href{https://doi.org/10.1088/1402-4896/aca2f7}{Phys. Scr. \textbf{97}, 125305 (2022).}


% Exact solutions of the Schrödinger equation with Dunkl derivative for the free-particle spherical waves, the pseudo-harmonic oscillator and the Mie-type potential
\bibitem{51} R. D. Mota, D. Ojeda-Guillén, \href{https://doi.org/10.1142/S0217732322500067}{Mod. Phys. Lett. A \textbf{37}, 2250006 (2022).}

% Effect of the two-parameter generalized Dunkl derivative on the two-dimensional Schrödinger equation
\bibitem{50} R. D. Mota, D. Ojeda-Guillén, \href{https://doi.org/10.1142/S0217732322502248}{Mod. Phys. Lett. A \textbf{37}, 2250224 (2022).} 

% Exact solution to two dimensional Dunkl harmonic oscillator in the Non-Commutative phase-space
\bibitem{46} S. Hassanabadi,  P. Sedaghatnia, W. S. Chung, B. C.  L\"{u}tf\"{u}o\u{g}lu, J. K\u{r}\'i\u{z},  H. Hassanabadi, \href{ https://doi.org/10.1140/epjp/s13360-023-03933-2}{Eur. Phys. J. Plus \textbf{138}, 331 (2023).}


% On the gauge invariance of Wigner–Dunkl quantum mechanics in the presence of a constant magnetic field
\bibitem{Junker2023} G. Junker, S. H. Dong, P. Sedaghatnia, W. S. Chung, H. Hassanabadi, \href{https://doi.org/10.1016/j.aop.2023.169336}{Ann. Phys. \textbf{454}, 169336 (2023).}

% Quasi-exactly solvable potentials in Wigner-Dunkl quantum mechanics
\bibitem{Quesne2024} C. Quesne, \href{https://doi.org/10.1209/0295-5075/ad2947}{EPL \textbf{145}, 62001 (2024).}

% On the path integral formulation of Wigner–Dunkl quantum mechanics
\bibitem{Junker2024} G. Junker, \href{https://doi.org/10.1088/1751-8121/ad213d}{J. Phys. A \textbf{57}, 075201 (2024).}

% One-dimensional Dunkl quantum mechanics: a path integral approach
\bibitem{66} A. Benchikha, B. Hamil, B. C. L\"{u}tf\"{u}o\u{g}lu, B. Khantoul, \href{https://doi.org/10.1088/1402-4896/ad7aaf}{Phys. Scr. \textbf{99}, 105274 (2024).}

%Path integral formulation for Dunkl-Dirac oscillator in (1+1)-dimensional space-time coordinates
\bibitem{Benzair20241} H. Benzair, T. Boudjedaa, M. Merad, \href{https://doi.org/10.1088/1402-4896/ad39b7}{Phys. Scr. \textbf{99}, 055261 (2024).}



%Path integral formalism for the Klein–Gordon-Dunkl oscillator in (1+d)-dimensional space-time coordinates
\bibitem{Benzair20242} H. Benzair, T. Boudjedaa, M. Merad, \href{https://doi.org/10.1088/1402-4896/ad6e2f}{Phys. Scr. \textbf{99}, 095256 (2024).}






% Dunkl-Schrödinger Equation with Time-Dependent Harmonic Oscillator Potential
\bibitem{67} A. Benchikha, B. Khantoul, B. Hamil, B. C. L\"{u}tf\"{u}o\u{g}lu, \href{https://doi.org/10.1007/s10773-024-05786-6}{Int. J. Theor. Phys. \textbf{63}, 248 (2024).}


% Closed-form solutions of the Dunkl–Klein–Gordon equation for two inverse power-law interactions
\bibitem{Schulze20245} A. Schulze-Halberg, \href{https://doi.org/10.1142/S0217732323501134}{Int. J. Mod. Phys. A \textbf{39}, 2450013 (2024).}

% Approximate bound state solutions of the Dunkl-Schrödinger equation for a hyperbolic double-well interaction
\bibitem{Schulze20243} A. Schulze-Halberg, \href{https://doi.org/10.1088/1402-4896/ad4fc8}{Phys. Scr. \textbf{99}, 075212 (2024).}

% Approximate Solutions of the Dunkl–Schrödinger Equation for the Hyperbolic Pöschl–Teller Potential
\bibitem{Schulze20244} A. Schulze-Halberg, \href{https://doi.org/10.1007/s00601-024-01931-3}{Few-Body Syst. \textbf{65}, 58 (2024).}

% The generalized Fokker–Planck equation in terms of Dunkl-type derivatives
\bibitem{Mota20241} R. D. Mota, D. Ojeda-Guill\'{e}n,  M. A. Xicot\'encatl, \href{https://doi.org/10.1016/j.physa.2024.129525}{Physica A \textbf{635},  129525 (2024).}

% The Dunkl–Fokker–Planck Equation in (1+1) Dimensions
\bibitem{Mota20242} R. D. Mota, D. Ojeda-Guill\'{e}n,  M. A. Xicot\'encatl,  \href{https://doi.org/10.1007/s00601-024-01898-1}{Few-Body Syst. \textbf{65}, 25 (2024).}


% Time-Dependent Dunkl-Schrödinger Equation with an Angular-Dependent Potential
\bibitem{67a} B. Khantoul, B. Hamil, A. Benchikha, B. C. L\"{u}tf\"{u}o\u{g}lu,  \href{https://doi.org/10.48550/arXiv.2408.02021}{arXiv:2408.02021 (2024).}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%








% Dunkl–Pauli equation in the presence of a magnetic field
\bibitem{Bouguerne20241} H. Bouguerne, B. Hamil, B. C. L\"{u}tf\"{u}o\u{g}lu, M. Merad, \href{https://doi.org/10.1007/s12648-024-03170-y}{Indian J. Phys. \textbf{98}, 4093 (2024).}






%
%% Superstatistics of the %Dunkl oscillator
%\bibitem{58} H. Hassanabadi, M. de Montigny, W. S. Chung, P. Sedaghatnia, \href{https://doi.org/10.1016/j.physa.2021.126154}{Physica A \textbf{580}, 126154 (2021).} 






%
%%  Ideal Bose gas and %blackbody radiation in the %Dunkl formalism
%\bibitem{60} F. Merabtine, B. Hamil, B. C. L\"{u}tf\"{u}o\u{g}lu, A. Hocine, M. Benarous, \href{https://doi.org/10.1088/1742-5468/acd106}{J. Stat. Mech. \textbf{2023}, 053102 (2023).}

%
%% Relativistic particles in %curved space with a static %metric using the Dunkl %derivative in the Dirac %operator
%\bibitem{65} P. Sedaghatnia, H. Hassanabadi, A. D. Alhaidari, W. S. Chung, \href{https://doi.org/10.1142/S0217751X22502232}{Int. J. Mod. Phys. A \textbf{37}, 2250223 (2023).}



%
%% The Dunkl oscillator on a %space of nonconstant %curvature: An exactly %solvable quantum model with %reflections
%\bibitem{45} A. Ballesteros, A. Najafizade, H. Panahi, H. Hassanabadi, S. H. Dong, \href{https://doi.org/10.1016/j.aop.2023.169543}{Ann. Phys. \textbf{460}, 169543 (2024).}

%
%% On Dunkl-Bose-Einstein Condensation in Harmonic Traps
%\bibitem{62} A. Hocine, B. Hamil, F. Merabtine, B. C. L\"{u}tf\"{u}o\u{g}lu, M. Benarous, \href{https://doi.org/10.31349/RevMexFis.70.051701}{Rev. Mex. Fís. \textbf{70},  051701 (2024).}


%
%% Condensation of ideal Dunkl–Bose gas in power-law traps
%\bibitem{63} A. Hocine, F. Merabtine, B. Hamil, B. C. L\"{u}tf\"{u}o\u{g}lu,  M. Benarous, \href{https://doi.org/10.1007/s12648-024-03311-3}{Indian J. Phys. \emph{early access} (2024).}



%%%%%%%%%%%%%%%%%%%%%%%%%%
%
\bibitem{68} M. A. Markov , V. V. Dodonov, V. I. Manko (Eds.), \emph{Invariants and the Evolution of Nonstationary Quantum Systems} (New-York: Nova Science Publishers,  1989).

% The quantum damped harmonic oscillator
\bibitem{69} C. I. Um, K. H. Yeon, T. F. George, \href{https://doi.org/10.1016/S0370-1573(01)00077-1}{Phys. Rep. \textbf{362}, 63 (2002).} 




% Linear and quadratic invariants for the transformed Tavis–Cummings model
\bibitem{70} M. S. Abdalla, P. G. L. Leach, \href{https://doi.org/10.1088/0305-4470/36/49/005}{ J. Phys. A: Math. Gen. \textbf{36}, 12205 (2003).}

% Wigner functions for time-dependent coupled linear oscillators via linear and quadratic invariant processes
\bibitem{71} M. S. Abdalla, P. G. L. Leach, \href{https://doi.org/10.1088/0305-4470/38/4/008}{J. Phys. A: Math. Gen. \textbf{38}, 881 (2005).}

% Time-Dependent 2D Harmonic Oscillator in Presence of the Aharanov-Bohm Effect
\bibitem{72} Y. Bouguerra, M. Maamache, A. Bounames, \href{https://doi.org/10.1007/s10773-006-9145-9}{Int. J. Theor. Phys. \textbf{45}, 1791 (2006).}

% Exact wavefunctions for a time-dependent Coulomb potential
\bibitem{73} S. Menouar, M. Maamache, Y. Saadi, J. R. Choi, \href{https://doi.org/10.1088/1751-8113/41/21/215303}{J. Phys. A: Math. Theor. \textbf{41}, 215303 (2008).}

% On the invariant method for the time-dependent non-Hermitian Hamiltonians
\bibitem{74} B. Khantoul, A. Bounames, M. Maamache, \href{https://doi.org/10.1140/epjp/i2017-11524-7}{Eur. Phys. J. Plus \textbf{132}, 258 (2017).}

% EXACT SOLUTIONS FOR TIME-DEPENDENT COMPLEXSYMMETRIC POTENTIAL WELL
\bibitem{75} B. Khantoul, A. Bounames, \href{https://doi.org/10.14311/AP.2023.63.0132}{Acta Polytech. \textbf{63}, 132 (2023).} 


% The Schrodinger and Pauli-Dirac oscillators in Noncommutative Phase Space
\bibitem{Santos2011} E. S. Santos, G. R. de Melo, \href{https://doi.org/10.1007/s10773-010-0529-5}{Int. J. Theor. Phys. \textbf{50}, 332 (2011).}


% Pauli oscillator in noncommutative space
\bibitem{Heddar2021} M. Heddar, M. Falek, M. Moumni, B. C. L\"{u}tf\"{u}o\u{g}lu, \href{https://doi.org/10.1142/S0217732321502801}{Mod. Phys. Lett. A \textbf{36}, 2150280 (2021).} 

% Exact solutions of  Pauli oscillator with Snyder-de Sitter algebra
\bibitem{Guerira2022} B. Guerira, M. Falek, M. Moumni, N. Belghar, \href{https://doi.org/10.1142/S0217751X22500695}{Int. J. Mod. Phys. A \textbf{37}, 2250069 (2022).}

% An Exact Quantum Theory of the Time‐Dependent Harmonic Oscillator and of a Charged Particle in a Time‐Dependent Electromagnetic Field 
\bibitem{Riesenfeld1969} H. R. Lewis, W. B. Riesenfeld, \href{https://doi.org/10.1063/1.1664991} {J. Math. Phys. \textbf{1010}, 1458 (1969).}

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{eurosym}
\usepackage{graphicx}
\usepackage[utf8]{inputenc}
\usepackage[T1]{fontenc}
\usepackage{indentfirst}
\usepackage[margin=0.6in,nomarginpar]{geometry}
\usepackage[final]{hyperref}
\usepackage{amsmath}
\usepackage{hyperref}
\usepackage{cite}
\usepackage{subcaption}
\usepackage{caption}
\usepackage{amssymb}
\usepackage{multirow}
\usepackage[table]{xcolor}
\usepackage{orcidlink}

\hypersetup{
colorlinks=true,
linkcolor=blue,
citecolor=blue,
filecolor=magenta,
urlcolor=blue
}

\begin{document}

\title{Time-dependent Dunkl-Pauli Oscillator}
\author{A. Benchikha \orcidlink{0009-0003-9848-0822}\thanks{%
benchikha4@yahoo.fr} \\
$^{1}$D\'{e}partement de EC, Facult\'{e} de SNV, Universit\'{e} Constantine 1 Fr\`{e}res
Mentouri,\\
Constantine, Algeria.\\
$^{2}$Laboratoire de Physique Math\'{e}matique et Subatomique,\\
LPMS, Universit\'{e} Constantine 1 Fr\`{e}res  Mentouri, Constantine,
Algeria. \and  B. Hamil \orcidlink{0000-0002-7043-6104} \thanks{%
hamilbilel@gmail.com} \\
Laboratoire de Physique Math\'{e}matique et Subatomique,\\
LPMS, Universit\'{e} Constantine 1 Fr\`{e}res Mentouri, Constantine,
Algeria.  \and B. C. L\"{u}tf\"{u}o\u{g}lu \orcidlink{0000-0001-6467-5005}\thanks{%
bekir.lutfuoglu@uhk.cz (Corresponding author)}, \\
Department of Physics, Faculty of Science, University of Hradec Kralove,\\
Rokitanskeho 62/26, Hradec Kralove, 500 03, Czech Republic.}
\date{\today }
\maketitle

\begin{abstract}
This study explores the time-dependent Dunkl-Pauli oscillator in two dimensions. We constructed the Dunkl-Pauli Hamiltonian, which incorporates a time-varying magnetic field and a harmonic oscillator characterized by time-dependent mass and frequency, initially in Cartesian coordinates. Subsequently, we reformulated the Hamiltonian in polar coordinates and analyzed the eigenvalues and eigenfunctions of the Dunkl angular operator, deriving exact solutions using the Lewis-Riesenfeld invariant method. Our findings regarding the total quantum phase factor and wave functions reveal the significant impact of Dunkl operators on quantum systems, providing precise expressions for wave functions and energy eigenvalues. This work enhances the understanding of quantum systems with deformed symmetries and suggests avenues for future research in quantum mechanics and mathematical physics.
\end{abstract}



\section{Introduction}
The Pauli equation, developed by Wolfgang Pauli as an extension of the Schr\"{o}dinger equation, integrates electron spin and magnetic field interactions to describe the dynamics of non-relativistic spin-1/2 particles, such as electrons, in electromagnetic fields \cite{1,2,3,4,5,6,7,8,9}. This equation has had a profound impact on quantum chemistry and particle physics, aiding in the classification of Lie symmetries for charged particles and quasi-relativistic Schr\"{o}dinger equations \cite{10,11,12,13}. It has also facilitated the discovery of new superintegrable systems exhibiting spin-orbit coupling, laying the theoretical foundation for atomic structure and the periodic table through quantum spin theory and the Pauli exclusion principle \cite{13,14,15,16,17,18,19,20,22}. Additionally, it supports Bohr's model of the hydrogen atom, which was previously regarded as empirical \cite{23,24}. Accurate solutions to the Pauli equation are crucial for predictive quantum chemistry, enabling precise predictions of multi-electron atoms and molecules \cite{25, 26, 27, 28}.

Dunkl operators, defined as commuting differential-difference operators linked to finite reflection groups \cite{Dunkl1989}, serve as a generalization of partial derivatives, enabling a broader framework for analyzing functions and their symmetries in pure mathematics \cite{Dunkl2008, Dunkl2014} and theoretical physics \cite{Kakei1996, Lapointe1996, Hikami1996}. Notably, the Dunkl-like operators first appeared in physics in 1951 through Yang's seminal work \cite{Yang1951}, drawing inspiration from the pioneering contributions of Wigner from the previous year \cite{Wigner1950}. Yang's work opened new frameworks in quantum field theory by proposing a generalized field quantization \cite{Green1953}, which significantly aided in the introduction of color degrees of freedom and the development of quantum chromodynamics \cite{Greenberg1964}. Subsequently, reflection operator deformed algebras have been employed in various subareas of physics, including particle physics and statistical mechanics, further enhancing our understanding of symmetries and interactions in these fields \cite{Plyushchay1994, Plyushchay1996, Plyushchay1997, Gamboa1999,   Plyushchay2000, Klishevich2001, Horvathy2004, Rodrigues2009, Horvathy2010, Ubriaco2014, Luo2020}.

In the past decade, Dunkl operators have increasingly been employed as substitutes for ordinary partial derivatives, leading to a richer algebraic structure in various quantum systems, both relativistic and nonrelativistic. This substitution not only enhances the understanding of solvability and integrability but also introduces new symmetries, providing deeper insights into the spectral properties and dynamics of these systems \cite{Genest20131, Genest20132, Genest20133, Genest20141, Genest20142, Vincent3, Isaac2016, Salazar2017, Salazar2018, Mota2019, Ghazouani2019, Ghazouani2020, Ghazouani2021, Hamil20221, Hamil20222, Hamil20223, Samira, 43, 46, 50, 51, Junker2023, Quesne2024, 66, 67, 67a, Schulze20243, Schulze20244, Schulze20245, Mota20241, Mota20242, Benzair20241, Benzair20242, Junker2024}. Recently, two authors of the present manuscript have investigated the Pauli equation within the Dunkl formalism in the presence of a magnetic field, deriving the Dunkl-Pauli Hamiltonian, which incorporates terms involving the Dunkl-Laplacian and additional potentials for magnetic field and spin interactions, with the solutions successfully presented in \cite{Bouguerne20241}. However, that was the stationary scenario, and in the non-stationary case, the dynamics are expected to reveal further complexities, which remain an open area of investigation. To address time-dependent scenarios, one can extend our previous analysis by considering a magnetic field that varies with time, along with a harmonic oscillator featuring both time-dependent mass and frequency potential. Such an extension results in a modified Hamiltonian that encapsulates the dynamic interactions between these variables. From there, one can express the wave function in polar coordinates and introduce the Dunkl-angular operator to facilitate solving the Hamiltonian. The eigenfunctions and eigenvalues of the Dunkl-angular operator are expected to play a central role in constructing solutions to the Dunkl-Pauli equation. Consequently, one can investigate cases with different eigenvalues of the reflection operators and derive the corresponding radial wave functions. Additionally, one can employ the invariant method to identify the time-dependent invariant and use unitary transformations to simplify the eigenvalue problem \cite{68, 69, 70, 71, 72, 73, 74, 75}.

The Pauli equation and the Pauli oscillator are both concepts in quantum mechanics that describe different behaviors of a spin-1/2 particle. The Pauli equation encompasses the general motion of the particle in electromagnetic fields. In contrast, the Pauli oscillator specifically pertains to the dynamics of such particles within a harmonic oscillator, incorporating spin and a magnetic field in a quadratic potential. Notably, the solution of the Pauli oscillator in conjunction with the Snyder-de Sitter algebra and noncommutative space yields intriguing predictions that further our understanding of quantum systems \cite{Santos2011, Heddar2021, Guerira2022}. Motivated by the facts summarized above, in this study, we aim to explain the dynamics of quantum systems described by the time-dependent Pauli oscillator within the framework of Dunkl formalism. We believe that the solution will provide new insights into the interplay between symmetry, algebra, and quantum dynamics, paving the way for potential applications in studying complex quantum systems with enhanced symmetry properties.

The paper is organized as follows: In Sec. \ref{sec2}, we derive the two-dimensional time-dependent Dunkl-Pauli Hamiltonian, first in Cartesian coordinates and then in polar coordinates. In Sec. \ref{sec3}, we employ the Lewis-Riesenfeld invariant method to investigate the wave function solutions. Subsequently, in Sec. \ref{sec4}, we discuss the total quantum phase factor and present the exact form of the wave functions. Finally, we conclude the manuscript with a brief conclusion section.

\section{Time-dependent Dunkl-Pauli oscillator in two dimensions} \label{sec2}

We consider the time-dependent Pauli equation with a harmonic oscillator potential in two dimensions, which represents the non-relativistic limit of the Dirac equation in standard quantum mechanics:
\begin{equation}
\left[ \frac{1}{2m\left( t\right) }\left( \overrightarrow{\sigma _{j}} \cdot \overrightarrow{\pi _{j}}\right) ^{2}+\frac{1}{2}m\left( t\right) \omega \left( t\right) r^{2}\right] \psi \left( \overrightarrow{r},t\right) =i\frac{\partial }{\partial t}\psi \left( \overrightarrow{r},t\right) ,  \label{1}
\end{equation}
where $m\left( t\right) $ denotes the time-dependent mass, while $\omega \left(t\right) $ corresponds to the time-dependent frequency of the harmonic oscillator. Here, $\sigma _{j}$ refers to the Pauli matrices and $\overrightarrow{\pi }$ are the kinetic momentum terms given by: 
\begin{equation}
\overrightarrow{\pi }=\overrightarrow{p}-\frac{e}{c}\overrightarrow{A}=\left( p_{1}-\frac{e}{c}A_{1},\text{ }p_{2}-\frac{e}{c}A_{2}\right) , \label{2}
\end{equation}%
with charge $e$ and speed of light $c$. In this case, the spinor wave function has two components:
\begin{equation}
\psi =\binom{\psi _{1}}{\psi _{2}}.  \label{3}
\end{equation}
We then adopt the Dunkl formalism by replacing the ordinary momentum operator with the Dunkl momentum operator ($\hbar=1$):
\begin{equation}
p_{j}=\frac{1}{i}D_{j},  \label{4}
\end{equation}%
where the Dunkl derivatives are defined as:%
\begin{equation}
D_{j}=\frac{\partial }{\partial x_{j}}+\frac{\nu _{j}}{x_{j}}\left(1-R_{j}\right) , \qquad {\text{for}} \qquad  j=1,2.   \label{5}
\end{equation}%
In this context, $\nu_j$ refers to the Wigner deformation parameters, while $R_j$ represents the reflection operators. To ensure consistency, the Wigner parameters must satisfy $\nu _{j} > -\frac{1}{2}$, and the reflection operators exhibit the following properties:
\begin{equation}
R_{j}f\left( x_{j}\right) =f\left( -x_{j}\right), \qquad R_{j}x_{i}=-\delta _{ij}x_{j}R_{j}, \qquad \text{and} \qquad R_{j}R_{j}=R_{j}R_{j} .   \label{7}
\end{equation}%
Employing the Dunkl second derivative 
\begin{equation}
D_{j}^{2}=\frac{\partial ^{2}}{\partial x_{j}^{2}}+\frac{2\nu _{j}}{x_{j}}\frac{\partial }{\partial x_{j}}-\frac{\nu _{j}}{x_{j}^{2}}\left(1-R_{j}\right) ,  \label{6}
\end{equation}%
we express the time-dependent Dunkl-Pauli Hamiltonian in Cartesian coordinates as follows:
\begin{equation}
H=\frac{1}{2m\left( t\right) }\bigg[ \pi _{1}^{2}+\pi _{2}^{2}+\sigma _{3}\Big[ \pi _{1},\pi _{2}\Big] \bigg] +\frac{1}{2}m\left( t\right) \omega \left( t\right) r^{2}.  \label{9}
\end{equation}%
We then choose the vector potential in the symmetric gauge, given by:
\begin{equation}
e\overrightarrow{A}=\frac{eB\left( t\right) }{2c}\left( -y\hat{i}+x\hat{j}\right) ,  \label{10}
\end{equation}%
where $\left(\hat{i},\hat{j}\right) $ represents the Cartesian orthonormal basis. By applying the Dunkl-deformed algebra and the definitions provided above, we derive the time-dependent Dunkl-Pauli Hamiltonian in the following form \cite{Bouguerne20241}:
\begin{equation}
H=-\frac{1}{2m\left( t\right) }\triangle _{D}+\frac{m\left( t\right) \Omega^{2}\left( t\right) }{2}\left( x^{2}+y^{2}\right) +i\frac{\omega _{c}\left(t\right) }{2}\left( xD_{2}-yD_{1}\right) -\frac{e}{2m\left( t\right) c}\left( 1+\nu _{1}R_{1}+\nu _{2}R_{2}\right) \sigma _{z}B\left( t\right), \label{11}
\end{equation}%
where the Dunkl-Laplacian is given by: 
\begin{equation}
\triangle _{D}=\frac{\partial ^{2}}{\partial x^{2}}+\frac{\partial ^{2}}{\partial y^{2}}+\frac{2\nu _{1}}{x}\frac{\partial }{\partial x}+\frac{2\nu_{2}}{y}\frac{\partial }{\partial y}-\frac{\nu _{1}}{x^{2}}\left(1-R_{1}\right) -\frac{\nu _{2}}{y^{2}}\left( 1-R_{2}\right).  \label{12}
\end{equation}%
Here, $\omega _{c}\left( t\right) =\frac{eB\left( t\right) }{m(t)c}$ is the time-dependent cyclotron frequency.

\subsection{Solutions in polar coordinates}
At this stage, we prefer to adopt polar coordinates to take advantage of the symmetries, as is customary. Defining the Cartesian coordinates in terms of polar coordinates as 
\begin{equation}
x=r\cos \theta \qquad \text{and} \qquad y=r\sin \theta ,  \label{13}
\end{equation}%
the time-dependent Dunkl-Pauli Hamiltonian takes the following form: 
\begin{align}
H\left( t\right) & =-\frac{1}{2m\left( t\right) }\frac{\partial ^{2}}{\partial r^{2}}-\frac{1+2\nu _{1}+2\nu _{2}}{2m\left( t\right) r}\frac{\partial }{\partial r}+\frac{m\left( t\right) }{2}\Omega ^{2}\left( t\right)r^{2}  \notag \\
& +\frac{\mathcal{J}_{\theta }^{2}-2\nu _{1}\nu _{2}\left(1-R_{1}R_{2}\right) }{2m\left( t\right) r^{2}}+\frac{\omega _{c}\left(t\right) }{2}\Big[ \mathcal{J}_{\theta }-g_{s}\left( 1+\nu _{1}R_{1}+\nu_{2}R_{2}\right) .\mathbf{S}_{z}\Big] ,  \label{14}
\end{align}%
where 
\begin{equation}
\Omega ^{2}\left( t\right) =\omega \left( t\right) ^{2}+\frac{\omega_{c}\left( t\right) ^{2}}{4},  \label{15}
\end{equation}
is the modified harmonic oscillator frequency. Here, $\mathbf{S}_{z}=\frac{\sigma _{z}}{2}$ represents the spin-$\frac{1}{2}$ operator, and $g_{s}$ is the free electron $g$-factor, with $g_{s}=2.0023 $. The Dunkl angular operator, denoted by $\mathcal{J}_{\theta }$, is defined as follows \cite{Genest20131, Genest20132, Genest20141, Genest20142}: 
\begin{equation}
\mathcal{J}_{\theta }=i\left( \frac{\partial }{\partial \theta }+\nu_{2}\cot \theta \left( 1-R_{2}\right) -\nu _{1}\tan \theta \left(1-R_{1}\right) \right) .  \label{17}
\end{equation}%
We then assume the time-dependent wave function takes the form
\begin{equation}
\psi _{m_{s}}\left( r,\theta ,t\right) =\phi \left( r,\theta ,t\right) \chi_{m_{s}},  \label{18}
\end{equation}%
where $\chi _{m_{s}}$ represents the spin function,%
\begin{equation}
\mathbf{S}_{z}\chi _{ms}=\frac{m_{s}}{2}\chi _{m_{s}} \qquad \text{for} \qquad
m_{s}=\pm 1,   \label{19a}
\end{equation}
and 
\begin{equation}
   \chi _{+1}=\binom{1}{0}, \qquad \chi _{-1}=%
\binom{0}{1}.  \label{19b}
\end{equation}
Subsequently, the time-dependent Dunkl-Pauli oscillator equation is expressed as:
\begin{align}
& \left[ -\frac{1}{2m\left( t\right) }\frac{\partial ^{2}}{\partial r^{2}}-\frac{1+2\nu _{1}+2\nu _{2}}{2m\left( t\right) r}\frac{\partial }{\partial r}+\frac{m\left( t\right) }{2}\Omega ^{2}\left( t\right) r^{2}+\frac{\mathcal{J}_{\theta }^{2}-2\nu _{1}\nu _{2}\left( 1-R_{1}R_{2}\right) }{2m\left(t\right) r^{2}}\right.  \notag \\
& \left. +\frac{\omega _{c}\left( t\right) }{2}\Big( \mathcal{J}_{\theta}-m_{s}\left( 1+\nu _{1}R_{1}+\nu _{2}R_{2}\right) \Big) \right] \phi \left( r,\theta ,t\right) =i\frac{\partial }{\partial t}\phi \left( r, \theta, t\right) .  \label{108}
\end{align}%
Here, one can perform the following transformation
\begin{equation}
\phi \left( r,\theta ,t\right) =e^{-\frac{i}{2}\left( \mathcal{J}_{\theta}-m_{s}\left( 1+\nu _{1}R_{1}+\nu _{2}R_{2}\right) \right) \int\limits^{t}\omega _{c}\left( t^{\prime }\right) dt^{\prime }}\mathcal{\digamma }(r,\theta ,t),  \label{109}
\end{equation}
to eliminate the last time-dependent term on the left-hand side of Eq. (\ref{108}), given by  $\frac{\omega_{c}\left( t\right) }{2}\Big( \mathcal{J}_{\theta }-m_{s}\left(1+\nu_{1}R_{1}+\nu _{2}R_{2}\right) \Big)$. Using this transformation, Eq. (\ref{108}) becomes:
\begin{equation}
\left\{ \frac{\mathbf{P}^{2}}{2m\left( t\right) }+\frac{\delta \left( \delta -1\right) -2\nu _{1}\nu _{2}\left( 1-R_{1}R_{2}\right) }{2m\left( t\right)r^{2}}+\frac{m\left( t\right) }{2}\Omega ^{2}\left( t\right) r^{2}\right\}\mathcal{\digamma }(r,\theta ,t)=i\frac{\partial }{\partial t}\mathcal{\digamma }(r,\theta ,t),  \label{a11}
\end{equation}%
where
\begin{eqnarray}
\mathbf{P}^{2} &=&\mathcal{P}^{2}+\frac{\mathcal{J}_{\theta }^{2}}{r^{2}}, \label{111} \\
\mathcal{P} &=&\frac{\hbar }{i}\left[ \frac{\partial }{\partial r}+\frac{\delta }{r}\right] ,  \label{112} \\
\delta &=&\nu _{1}+\nu _{2}+\frac{1}{2},  \label{113}
\end{eqnarray}%
with $\left[ r,\mathcal{P}\right] =i\hbar .$ We observe that the time-dependent Dunkl-Pauli equation mimics the time-dependent
Dunkl-Shr\"odinger equation, as outlined in equation (9) in   \cite{67}.

\section{Lewis-Riesenfeld invariant approach} \label{sec3}

To obtain an exact solution, we utilize the Lewis-Riesenfeld invariant approach  \cite{Riesenfeld1969}, requiring that the Hamiltonian of the system and the invariant $I(t)$ satisfy the Lewis-Riesenfeld invariant equation:
\begin{equation}
\frac{dI(t)}{dt}=\frac{\partial I(t)}{\partial t}+\frac{1}{i}[I(t),\tilde{H}(t)]=0,  \label{114}
\end{equation}%
where the Hamiltonian is given by on the left-side of Eq. \eqref{a11}
\begin{equation}
\tilde{H}(t)=\frac{\mathbf{P}^{2}}{2m\left( t\right) }+\frac{\delta \left(\delta -1\right) -2\nu _{1}\nu _{2}\left( 1-R_{1}R_{2}\right) }{2m\left(t\right) r^{2}}+\frac{m\left( t\right) }{2}\Omega ^{2}\left( t\right) r^{2}.
\label{new Hamiltonian}
\end{equation}
According to this approach the solution of Eq. (\ref{a11}) can be related to the solution of the eigenvalue problem of the invariant
\begin{equation}
I(t)\mathcal{F}(r,\theta ,t)=\varepsilon \mathcal{F}(r,\theta ,t),
\label{I1}
\end{equation}%
with the associated phase factor given by 
\begin{equation}
\mathcal{\digamma }(r,\theta ,t)=e^{i\eta (t)}\mathcal{F}(r,\theta ,t),
\label{115}
\end{equation}%
where the phase $\eta (t)$ can be determined from the equation 
\begin{equation}
\frac{d}{dt}\eta (t)=\Big{\langle} \mathcal{F}(r,\theta ,t)\Big{|}i\frac{\partial }{\partial t}-\tilde{H}\Big{|}\mathcal{F}(r,\theta ,t)\Big{\rangle} .  \label{PH}
\end{equation}%
To find the exact invariants for the system, we define three generators $T_{1}$, $T_{2}$ and $T_{3}$ from the Lie algebra of the group SL(2,R), which satisfy the following commutation relations: 
\begin{equation}
\lbrack T_{1},T_{2}]=-2i\hbar T_{3},\qquad \lbrack T_{2},T_{3}]= 4i\hbar T_{2},\qquad \lbrack T_{1},T_{3}]= - 4i\hbar T_{1}.  \label{37a}
\end{equation}%
and we form the invariant as follows: 
\begin{equation}
I(t)=\frac{1}{2}\Big(\alpha T_{1}+\beta T_{2}+\gamma T_{3}\Big).  \label{37b}
\end{equation}%
Subsequently, by applying Eq. (\ref{114}), we obtain a set of coupled equations whose solutions are expressed in terms of real functions of time, $\alpha$, $\beta$, and $\gamma $:
\begin{equation}
\left\{ \begin{aligned} \alpha &= \rho^{2}, \\ \beta &= \frac{1}{\rho^{2}} + m^{2}\dot{\rho}^{2}, \\ \gamma &= -m\rho \dot{\rho}. \end{aligned}\right. \label{37c}
\end{equation}%
Thus, the invariant can be explicitly represented as: 
\begin{equation}
I=\frac{1}{2}\left[ \left( \frac{1}{\rho ^{2}}+m^{2}\dot{\rho}^{2}\right) r^{2}+\rho ^{2}\left( \mathbf{P}^{2}+\frac{\delta \left( \delta -1\right)-2\nu _{1}\nu _{2}\left( 1-R_{1}R_{2}\right) }{r^{2}}\right) -\rho \dot{\rho}m\left( r\mathcal{P}+\mathcal{P}r\right) \right] ,  \label{38}
\end{equation}%
where $\rho $ obeys the Ermakov-Pinney equation, which is given by: 
\begin{equation}
\ddot{\rho}+\frac{\dot{m}}{m}\dot{\rho}+\Omega ^{2}\rho =\frac{1}{m^{2}\rho
^{3}}.  \label{38a}
\end{equation}%
To solve the eigenvalue equation presented in Eq. \eqref{I1}, where the invariant is given in Eq. \eqref{38}, we introduce the following unitary transformation%
\begin{equation}
\mathcal{F}(r,\theta ,t)=U\left( r\right) \mathcal{G}(r,\theta ),  \label{40}
\end{equation}%
with the unitary operator
\begin{equation}
U\left( r\right) =\exp \left( \frac{im\dot{\rho}}{2\rho }r^{2}\right).
\label{41}
\end{equation}%
This transformation leads to the the operator $I$ being modified to $I^{\prime }$, according to the relation $I^{\prime }$:
\begin{equation}
I^{\prime }=U^{+}IU=\frac{1}{2}\left[ \rho ^{2}\left( \mathbf{P}^{2}+\frac{\delta
\left( \delta -1\right) -2\nu _{1}\nu _{2}\left( 1-R_{1}R_{2}\right) }{r^{2}}%
\right) +\frac{1}{\rho ^{2}}r^{2}\right] .  \label{42}
\end{equation}%
Consequently, Eq. \eqref{I1} maps to
\begin{equation}
I^{^{\prime }}(t)\mathcal{G}(r,\theta )=\varepsilon \mathcal{G}(r,\theta ).
\label{43}
\end{equation}%
Assuming
\begin{equation}
    \mathcal{G}(r,\theta )=r^{-\delta }\mathcal{Q}(r)\Theta _{\epsilon
}\left( \theta \right) ,
\end{equation}
the wave equation given in Eq. (\ref{43}) separates into two differential equations:
\begin{equation}
\mathcal{J}_{\theta }\Theta _{\epsilon }\left( \theta \right) =\lambda
_{\epsilon }\Theta _{\epsilon }\left( \theta \right) ,  \label{angular}
\end{equation}
\begin{equation}
\left[ \frac{\partial ^{2}}{\partial \varkappa ^{2}}-\varkappa ^{2}-\frac{\left(
\sigma _{l}^{\epsilon }\right) ^{2}-\frac{1}{4}}{\varkappa ^{2}}%
+2\varepsilon _{n,l}^{\epsilon }\right] \mathcal{Q}(\varkappa )=0,
\label{radial1}
\end{equation}%
in terms of the new variable $\varkappa =\frac{r}{\rho }$, and 
\begin{equation}
\sigma _{l}^{\epsilon }=\sqrt{\lambda _{\epsilon }^{2}+\left( \nu
_{1}+\epsilon \nu _{2}\right) ^{2}}.  \label{116}
\end{equation}
Here, $\epsilon =\epsilon _{1}\epsilon _{2}=\pm 1$, where each $\epsilon _{i}=\pm1,$ represents the eigenvalues associated with the reflection operators $R_{i}$. We then determine the eigenfunctions, $\Theta _{\epsilon }\left( \theta \right) $ and the corresponding eigenvalues $\lambda _{\epsilon }$ for two different cases:

\begin{itemize}

\item \textbf{First case}, $\epsilon =1$\textbf{:} 

This case corresponds to $\epsilon _{1}=\epsilon _{2}=1$ or $\epsilon _{1}=\epsilon _{2}=-1$ subcases, and the solution for $\Theta _{+1}\left( \theta \right) $ are given in \cite{Bouguerne20241} as follows: 
\begin{equation}
\Theta _{+1}\left( \theta \right) =a_{l}\mathbf{P}_{l}^{\left( \nu
_{1}+1/2,\nu _{2}+1/2\right) }\left( -2\cos \theta \right) \pm \acute{a}%
_{l}\sin \theta \cos \theta \mathbf{P}_{l-1}^{\left( \nu _{1}+1/2,\nu
_{2}+1/2\right) }\left( -2\cos \theta \right) ,  \label{24}
\end{equation}%
where $\mathbf{P}_{l}^{\left( \alpha ,\beta \right) }$ are Jacobi
polynomials and the eigenvalue is:
\begin{equation}
\lambda _{+}=\pm 2\sqrt{l\left( l+\nu _{1}+\nu _{2}\right) },\qquad
l\in  \mathbb{N}^{\ast }.  \label{25}
\end{equation}

\item \textbf{Second case, }$\epsilon =-1$\textbf{:} 

This case involves two sub-cases: $\left( \epsilon _{1},\epsilon _{2}\right) =\left( +1,-1\right) $
or $\left( \epsilon _{1},\epsilon _{2}\right) =\left( -1,+1\right) .$ Here,
the angular eigenfunction $\Theta _{-1}\left( \theta \right) $ is given by:%
\begin{equation}
\Theta _{-1}\left( \theta \right) =b_{l}\mathbf{P}_{l-1/2}^{\left( \nu
_{1}+1/2,\nu _{2}-1/2\right) }\left( -2\cos \theta \right) \pm \acute{b}%
_{l}\sin \theta \mathbf{P}_{l-1/2}^{\left( \nu _{1}-1/2,\nu _{2}+1/2\right)
}\left( -2\cos \theta \right) ,  \label{26}
\end{equation}%
and the corresponding eigenvalue is: \begin{equation}
\lambda _{-}=\pm 2\sqrt{\left( l+\nu _{1}\right) \left( l+\nu _{2}\right) }%
,\ \text{\ \ }l\in \left\{ 1/2,3/2,5/2, \cdots\right\} .  \label{27}
\end{equation}%
\end{itemize}
Subsequently, the solution of Eq. (\ref{radial1}) is given by:
\begin{equation}
\mathcal{Q}(\varkappa )=\mathcal{C}_{r}\varkappa ^{\sigma _{l}^{\epsilon }+%
\frac{1}{2}}e^{-\frac{\varkappa ^{2}}{2}}L_{n}^{\sigma _{l}^{\epsilon
}}\left( \varkappa ^{2}\right) ,  \label{28}
\end{equation}%
where the eigenvalues of the invariant are
\begin{equation}
\varepsilon _{n,l}^{\epsilon }=2n+\sigma _{l}^{\epsilon }+1.\text{ \ \ \ \ \ 
}\ n=0,1, \cdots .  \label{46}
\end{equation}
Here, $\mathcal{C}_{r}$ represents the normalization constant. 

\newpage

\section{Total quantum phase and wave function} \label{sec4}

By applying a unitary transformation and utilizing the Ermakov-Pinney equation, we simplify the expression in Eq. \eqref{PH}, yielding:
\begin{equation}
\dot{\eta}\left( t\right) =-\frac{\varepsilon _{n,l}}{m\rho ^{2}}+\Big{\langle} \mathcal{F}(r,\theta ,t)\Big{\vert} i\frac{\partial }{\partial
t}-\frac{\dot{\rho}}{2\rho }\left( r\mathcal{P}+\mathcal{P}r\right)
\Big{\vert} \mathcal{F}(r,\theta ,t))\Big{\rangle}.  \label{diff phase}
\end{equation}%
where it is straightforward to show that:
\begin{equation}
\Big{\langle} \mathcal{F}(r,\theta ,t)\Big{\vert} i\frac{\partial }{\partial t%
}-\frac{\dot{\rho}}{2\rho }\left( r\mathcal{P}+\mathcal{P}r\right)
\Big{\vert} \mathcal{F}(r,\theta ,t)\Big{\rangle} =0,  \label{102}
\end{equation}%
Thus, the phase can be written as
\begin{equation}
\eta _{l,n}^{\epsilon }\left( t\right) =-\left( 2n+\sigma _{l}^{\epsilon
}+1\right) \int^{t}\frac{dt^{\prime }}{m\left( t^{\prime }\right) \rho
\left( t^{\prime }\right) ^{2}},  \label{105}
\end{equation}%
Combining these results, we find that the total time-dependent wave function for Eq. \eqref{1} takes the form:
\begin{eqnarray}
\psi _{n,l,m_{s}}^{\epsilon _{1},\epsilon _{2}}\left( r,\theta ,t\right) &=&%
\mathcal{C}_{r}\frac{r^{\sigma _{l}^{\epsilon }-\delta +\frac{1}{2}}}{\rho
^{\sigma _{l}^{\epsilon }+\frac{1}{2}}}\exp \left\{ \left( im\rho \dot{\rho}%
-1\right) \frac{r^{2}}{2\rho ^{2}}\right.  \notag \\
&&\left. +i\left[ \frac{1}{2}\left( \lambda _{\epsilon }-m_{s}\left( 1+\nu
_{1}\epsilon _{1}+\nu _{2}\epsilon _{2}\right) \right) \int\limits^{t}\omega
_{c}\left( t^{\prime }\right) dt^{\prime }-\left( 2n+\sigma _{l}^{\epsilon
}+1\right) \int^{t}\frac{dt^{\prime }}{m\left( t^{\prime }\right) \rho
\left( t^{\prime }\right) ^{2}}\right] \right\}  \notag \\
&&\times L_{n}^{\sigma _{l}^{\epsilon }}\left( \frac{r^{2}}{\rho ^{2}}%
\right) \Theta _{\epsilon }\left( \theta \right) \chi _{m_{s}}.
\label{total wave function}
\end{eqnarray}%
Using the total wave function and Eq. (\ref{109}), we obtain the total phase factor $\mu _{l,n}^{\epsilon }\left(t\right) $in the following form:
\begin{equation}
\mu _{l,n}^{\epsilon }\left( t\right) =\frac{1}{2}\left( \lambda _{\epsilon
}-m_{s}\left( 1+\nu _{1}\epsilon _{1}+\nu _{2}\epsilon _{2}\right) \right)
\int\limits^{t}\omega _{c}\left( t^{\prime }\right) dt^{\prime }-\left(
2n+\sigma _{l}^{\epsilon }+1\right) \int^{t}\frac{dt^{\prime }}{m\left(
t^{\prime }\right) \rho \left( t^{\prime }\right) ^{2}}.  \label{phase T}
\end{equation}%

\noindent We also note that the wave functions in Eq. (\ref{total wave function}) depend on the values of $\epsilon$, so the solutions have to be distinguished by two cases:

\begin{itemize}

\item  \textbf{First case}, $\epsilon =1$\textbf{:}

In this case,  we have $\sigma _{l}^{+}=\sqrt{\lambda
_{+}^{2}+\left( \nu _{1}+\nu _{2}\right) ^{2}}$ and $\lambda _{+}=\pm 2\sqrt{%
l\left( l+\nu _{1}+\nu _{2}\right) }.$

\begin{enumerate}
    \item For $\epsilon _{1}=1,$ $\epsilon _{2}=1,$ we get
\small
\begin{eqnarray}
\psi _{n,l,m_{s}}^{+,+}\left( r,\theta ,t\right) &=&\mathcal{C}_{r}\frac{%
r^{\sigma _{l}^{+}-\delta +\frac{1}{2}}}{\rho ^{\sigma _{l}^{+}+\frac{1}{2}}}%
\exp \left\{ \left( im\rho \dot{\rho}-1\right) \frac{r^{2}}{2\rho ^{2}}%
\right.  \notag \\
&+i&\left. \left[ \frac{1}{2}\left( \lambda _{+}-m_{s}\left( 1+\nu _{1}+\nu
_{2}\right) \right) \int\limits^{t}\omega _{c}\left( t^{\prime }\right)
dt^{\prime }-\left( 2n+\sigma _{l}^{+}+1\right) \int^{t}\frac{dt^{\prime }}{%
m\left( t^{\prime }\right) \rho \left( t^{\prime }\right) ^{2}}\right]
\right\}  \notag \\
&\times& L_{n}^{\sigma _{l}^{+}}\left( \frac{r^{2}}{\rho ^{2}}\right) \Theta
_{+1}\left( \theta \right) \chi _{m_{s}}.  \label{wave1}
\end{eqnarray}

    \item For $\epsilon _{1}=-1,\epsilon _{2}=-1,$ we get
\small
\begin{eqnarray}
\psi _{n,l,m_{s}}^{-,-}\left( r,\theta ,t\right) &=&\mathcal{C}_{r}\frac{%
r^{\sigma _{l}^{+}-\delta +\frac{1}{2}}}{\rho ^{\sigma _{l}^{+}+\frac{1}{2}}}%
\exp \left\{ \left( iM\rho \dot{\rho}-1\right) \frac{r^{2}}{2\rho ^{2}}%
\right.  \notag \\
&+i&\left. \left[ \frac{1}{2}\left( \lambda _{+}-m_{s}\left( 1-\nu _{1}-\nu
_{2}\right) \right) \int\limits^{t}\omega _{c}\left( t^{\prime }\right)
dt^{\prime }-\left( 2n+\sigma _{l}^{+}+1\right) \int_{0}^{t}\frac{dt^{\prime
}}{M\left( t^{\prime }\right) \rho \left( t^{\prime }\right) ^{2}}\right]
\right\}  \notag \\
&\times& L_{n}^{\sigma _{l}^{+}}\left( \frac{r^{2}}{\rho ^{2}}\right) \Theta
_{+1}\left( \theta \right) \chi _{m_{s}}.  \label{wave2}
\end{eqnarray}
\end{enumerate}

\item  \textbf{Second case}, $\epsilon =-1$\textbf{:} 

In this case, we have  $\sigma _{l}^{-}=\sqrt{\lambda_{-}^{2}+\left( \nu _{1}-\nu _{2}\right) ^{2}}$ and $\lambda _{-}=\pm 2\sqrt{%
\left( l+\nu _{1}\right) \left( l+\nu _{2}\right) }.$

\begin{enumerate}
    
    \item For $\epsilon _{1}=-1,$ $\epsilon _{2}=1,$ we have
\small    
\begin{eqnarray}
\psi _{n,l,m_{s}}^{-,+}\left( r,\theta ,t\right) &=&\mathcal{C}_{r}\frac{%
r^{\sigma _{l}^{-}-\delta +\frac{1}{2}}}{\rho ^{\sigma _{l}^{-}+\frac{1}{2}}}%
\exp \left\{ \left( im\rho \dot{\rho}-1\right) \frac{r^{2}}{2\rho ^{2}}%
\right.  \notag \\
&+i&\left. \left[ \frac{1}{2}\left( \lambda _{-}-m_{s}\left( 1-\nu _{1}+\nu
_{2}\right) \right) \int\limits^{t}\omega _{c}\left( t^{\prime }\right)
dt^{\prime }-\left( 2n+\sigma _{l}^{-}+1\right) \int^{t}\frac{dt^{\prime }}{%
m\left( t^{\prime }\right) \rho \left( t^{\prime }\right) ^{2}}\right]
\right\}  \notag \\
&\times& L_{n}^{\sigma _{l}^{-}}\left( \frac{r^{2}}{\rho ^{2}}\right) \Theta
_{-1}\left( \theta \right) \chi _{m_{s}}.  \label{wave 3}
\end{eqnarray}

    \item For $\epsilon _{1}=1,$ $\epsilon _{2}=-1,$ we have

\small
\begin{eqnarray}
\psi _{n,l,m_{s}}^{+,-}\left( r,\theta ,t\right) &=&\mathcal{C}_{r}\frac{%
r^{\sigma _{l}^{-}-\delta +\frac{1}{2}}}{\rho ^{\sigma _{l}^{-}+\frac{1}{2}}}%
\exp \left\{ \left( iM\rho \dot{\rho}-1\right) \frac{r^{2}}{2\rho ^{2}}%
\right.  \notag \\
&+i&\left. \left[ \frac{1}{2}\left( \lambda _{-}-m_{s}\left( 1+\nu _{1}-\nu
_{2}\right) \right) \int\limits^{t}\omega _{c}\left( t^{\prime }\right)
dt^{\prime }-\left( 2n+\sigma _{l}^{-}+1\right) \int^{t}\frac{dt^{\prime }}{%
M\left( t^{\prime }\right) \rho \left( t^{\prime }\right) ^{2}}\right]
\right\}  \notag \\
&\times& L_{n}^{\sigma _{l}^{-}}\left( \frac{r^{2}}{\rho ^{2}}\right) \Theta
_{-1}\left( \theta \right) \chi _{m_{s}}.  \label{wave4}
\end{eqnarray}

\end{enumerate}

\end{itemize}

\normalsize
For the time-independent case, where $\omega _{c}=C^{st}$ and $m=C^{st}$, we can take  $\omega
\left( t\right) \rightarrow 0$. Thus, we get
\begin{equation}
\Omega =\frac{\omega _{c}}{2},\text{ \ }\rho =\frac{1}{\sqrt{\frac{1}{2}%
\omega _{c}m}},  \label{constants}
\end{equation}
and the total phase $\mu _{l,n}^{\epsilon }\left( t\right) $ becomes the same with the energy
spectrum $E_{n,l}^{\epsilon }$ of the stationary Dunkl-Pauli oscillator equation
\begin{equation}
\mu _{l,n}^{\epsilon }\left( t\right) =-E_{n,l}^{\epsilon }t,
\label{phase t}
\end{equation}
with
\begin{equation}
E_{n,l}^{\epsilon }=\frac{\omega _{c}}{2}\Big[ 2n+\sigma _{l}^{\epsilon}+1-\lambda _{\epsilon }+m_{s}\left( 1+\nu _{1}\epsilon _{1}+\nu
_{2}\epsilon _{2}\right) \Big] .  \label{energy}
\end{equation}

\noindent This energy spectrum result depends on the eigenvalues $\epsilon_1$ and $\epsilon_2$ of the reflection operators $R_1$ and $R_2$, which aligns with the findings presented in reference \cite{Bouguerne20241}.

\section{Conclusion}

In this study, we investigated the time-dependent Dunkl-Pauli oscillator in two dimensions, utilizing the Pauli equation as our foundational framework. By introducing the Dunkl formalism, we successfully constructed the Dunkl-Pauli equation in Cartesian coordinates, incorporating reflection symmetries and deformed momentum operators. This formulation enabled us to derive the Dunkl-Pauli Hamiltonian, influenced by a time-varying magnetic field and a harmonic oscillator characterized by time-dependent mass and frequency. Subsequently, we expressed the modified Hamiltonian in polar coordinates. Through an analysis of the eigenvalues and eigenfunctions of the Dunkl angular operator, we obtained exact wave function solutions using the Lewis-Riesenfeld invariant method. Additionally, we derived the total quantum phase factors for various reflection cases. Our results underscored the distinctive impact of Dunkl operators on the quantum system, providing precise expressions for the wave functions and energy eigenvalues. These findings illustrate the potential of Dunkl operators to deepen our understanding of quantum systems with deformed symmetries, paving the way for future research in quantum mechanics and mathematical physics.


\section*{Acknowledgments}

B. C. L. is grateful to Excellence Project PřF UHK 2211/2023-2024 for the financial support.

\section*{Data Availability Statement} 

Data sharing is not applicable to this article as no new data were created or analyzed in this study.

\section*{Conflict of Interests} 

The authors have no conflicts to disclose.

\end{document}